OpenStudy (anonymous):

A line with y-intercept of 2 that is parallel to 2x + y = -5.

terenzreignz (terenzreignz):

If the line is parallel to 2x + y = -5, then that line and the line 2x+y=5 must have the same ___________? Fill int hat blank :D

terenzreignz (terenzreignz):

*Fill in that blank

OpenStudy (jackellyn):

Solve for y to put the equation in y=mx+b form to find out the ________ (what @terenzreignz is asking you to fill in the blank).